In 10 seconds, a disc accelerates uniformly from rest to a rational velocity of 20 rotations per second. Calculate the flywheel's angular acceleration in rad/s²?

1
Expert's answer
2021-11-30T17:44:49-0500

"\omega=\omega_0+\epsilon t\to \epsilon=(\omega-\omega_0)/t=(2\pi\nu-2\pi\nu_0)/t="


"=(2\cdot3.14\cdot20-0)/10=12.56\ (rad/s^2)" . Answer